The J.O. Supply Company buys calculators from a Korean supplier. The probability of a defective calculator is 10%. 100 calculators are selected at random. What is the expected number of defectives? What is the standard deviation? What is the probability that 3 or more calculators will be defective?
Answer by stanbon(75887)

The J.O. Supply Company buys calculators from a Korean supplier.
The probability of a defective calculator is 10%.
100 calculators are selected at random.
What is the expected number of defectives?
E(x) = np = 100*0.10 = 10
-----------------------------------
What is the standard deviation?
std = sqrt(npq) = sqrt(10*0.9) = 3
-----------------------------------
What is the probability that 3 or more calculators will be defective?
P(3 or more) = 1 - P(0<= x <=2) = 1 - binomcdf(100,0.1,2) = 0.9900
----------------------
Cheers,
Stan H.
======================